OpenStudy (anonymous):

MULTIPLY: n^2-3n-4 8n+24 --------- X --------- n+1 n^2-n-12

OpenStudy (anonymous):

=(n-4)(n+1) 8(n+3) --------- x ______ (n+1) (n-4)(n+3) =(n-4) x 8 ___ (n-4) =8